1. Mrs Smith has an order for a volume based enteral feeding. Her order reads: 1800 ml FS Osmolite 1.2 per day. The enteral feeding is started at 8am Friday. Mrs Smith is off her enteral feeding for 3 hours (1pm to 4pm) getting a CAT scan.

1A. What is the initial rate of the tube feeding?

2A. When Mrs Smith returns from her test, what rate is the enteral feeding run at until completion at 8am on Saturday?

  

2. Mr Smith has an order for a volume based enteral feeding. His order reads: 1500 ml FS Jevity 1.5 per day. The enteral feeding is started Monday at 6am. Mr Smith’s tube feedings are turned off from 12pm-2pm while the RN is waiting for a new bottle of tube feeding to arrive on the floor. His feeding is also held from 5pm to 7pm while having a new central line placed.

  2A. What is the initial rate of the tube feeding?   

2B. When Mr Smith’s enteral feedings are resumed at 7pm, what rate will the enteral feeding run at until completion at 6am Tuesday?       

Subject-Nutrition
Answer #1

1) a) Initially 1800 ml had to be infused in 24 hours.

So rate of infusion = 1800/24 = 75 ml/hour

b)From 8 AM to 1 PM , volume infused =75×5 =375 ml

Volume to be infused=1800-375 =1425 ml

Time left = 16 hours

Rate = 1425/16 = 89.06 ml/hour

2)a) Initial rate =1500/24 = 62.5 ml/hour

b) Total time =6 AM to 12 PM and 2-5 PM

Volume infused till 7 PM =62.5 *9 = 562.5 ml

Volume time be infused = 1500 -562.5 = 937.5 ml

Time = 11 hours (7 PM to 6 AM)

Rate = 937.5/11 = 85.23 ml/hour